I am stuck on correct answer for:
3 over 2 - 5 over 6 =
I believe answer is 2 over 3 but I need to verify my work. Can you show me the breakdown to get correct resolve?

Answer by anjulasahay(30)

I am stuck on correct answer for:
3 over 2 - 5 over 6 =
I believe answer is 2 over 3 but I need to verify my work. Can you show me the breakdown to get correct resolve?
ans:
as given (3/2)-(5/6)=(9-5)/6=4/6=2/3
your ans is correct.